26
May
2017
16:04

We are experiencing interruptions to our connectivity with one of our upstream providers, Pulsant, this afternoon.

As we use diverse suppliers for transit we switched the Pulsant transit off until they had fixed their issue.

We apologise for any inconvenience caused.

Posted by Yolanda Zyczynska

04
May
2017
14:25

Update: Pulsant have notified us that they will be performing emergency maintenance between 21:00 and 23:00 this evening related to the earlier outage.

This maintenance should be considered as an ‘at risk’ time, rather than a definite outage.

Earlier message:

We unfortunately experienced network outages outwith our network earlier today.

Our infrastructure was working normally but the issues were with one transit provider.

The issue was with Pulsant; more information can be found at http://status.pulsant.com/

We apologise for the three blips these issues caused.

Posted by Paul Malkowski

10
Apr
2017
16:25

We would like to offer our apologies for the brief network outage that happened in the last hour.

We were monitoring traffic using debug on one of the main switches and unfortunately ran out of memory.

As soon as we realised what was happening we reversed the debug monitoring and normal service was resumed.

Posted by Paul Malkowski

05
Apr
2017
14:20

We experienced a blip with Pulsant transit coming out of Pulsant 1 & 2 in Maidenhead early this afternoon.

As we use diverse suppliers for transit we switched the Pulsant transit off until they had fixed their issue.

They currently have a workaround in place and everything seems to be normal again.

We have switched our Pulsant transit back on.

Posted by Paul Malkowski

02
Mar
2017
13:27

We would like to apologise for not informing everyone of the maintenance window between 5pm and 7pm on Tuesday 28 February 2017.

We were moving a physical between racks which hosted some of our servers including knockando our mail server.

Posted by Paul Malkowski

09
Dec
2016
20:08

We would like to apologise for the outage affecting various services between approximately 18:15 and 18:45 this evening.

This was caused by a configuration error whilst deploying a private interconnection with a specific customer. Unfortunately a layer2 routing loop was created, which flooded a core network ethernet segment with broadcast traffic.

Posted by Kevin Hones

09
Dec
2016
14:32

There was an unexpected switch cluster (Beatrix) reboot.

This is now working again.

We will investigate the cause of this and post an update.

Posted by Paul Malkowski

03
Nov
2016
18:19

Shortly after 18:00 we suffered an interconnectivity glitch lasting a few minutes affecting leased line (fibre and EoFTTC) traffic.
All services are restored and we would like to apologise for the interruption in service.

Posted by Kevin Hones

25
Jul
2016
12:51

There was an outage around 12.30 today for around 10 minutes in one of our data centers.,

This provides connectivity for all leased lines and EoFTTC in our network.

This seems similar to the outage we experienced last week.

This has been fixed and we are currently looking into the exact cause of this.

Posted by Paul Malkowski

21
Jul
2016
10:56

The TalkTalk DSL connections have gone down again. As soon as we have had an update from our supplier we will post an update.

Update 10:58 – service to the lines has been restored again.

Posted by Yolanda Zyczynska

21
Jul
2016
08:18

All DSL connections from us using TalkTalk Business circuits are currently down. Our wholesale supplier is investigating – the fault is currently believed to be within TalKTalk’s network. We will publish more information as soon as we have it.

Update 08:50 – our supplier has notified us that the problem is a power issue in Telehouse North. It seems that many providers are severely affected by this, including TalkTalk and BT. There is no ETA yet, but we will pass on any updates as we receive them.

Update 08:55 – our supplier has stated “This has been confirmed as a power issue within Telehouse North affecting multiple areas. We will post a further update by 10:30 BST.”

Update 10:25 – service is being restored to the affected TalkTalk DSL lines. We will be observing the lines to to make sure all lines come back up.

Update 10:40 – our supplier has updated us with the following “The suite had power, but at the exact moment of the wider Telehouse power issue today the in-rack UPS unit appears to have failed. We believe a power surge may have damaged the equipment but engineers are looking through the logs to try and understand exactly what happened.”

Posted by Kevin Hones

20
Jul
2016
08:29

There was an outage around 8am this morning for 30 minutes in one of our data centers.,

This provides connectivity for all leased lines and EoFTTC in our network.

This has been fixed and we are currently looking into the exact cause of this.

Posted by Paul Malkowski

Current Status

    GoodNo current issues

Status Archive